Circular No. 6302

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

3C 279 F. Makino, H. Inoue, T. Kii, F. Nagase, and A. Yamashita, Institute

of Space and Astronautical Science, on behalf of the ASCA team, report: “Following the report of a γ-ray flare of 3C 279 (IAUC 6294), ASCA observed this region during Jan. 27–28. From a quick-look analysis, the 2- to 10-keV flux is estimated to be 1.2 × 10^{−11} erg cm^{−2} s^{−1}. The spectrum is consistent with a power law with a photon index of ∼ 1.7. According to the previous ASCA observations, the 2- to 10-keV flux was (6–9) × 10^{−12} erg cm^{−2} s^{−1} from the middle of 1993 to the beginning of 1995, and was (1.0–1.6) × 10^{−11} erg cm^{−2} s^{−1} in the middle of 1995. 3C 279 seems to have shown no large intensity change in the band 2–10 keV since the middle of 1995, in spite of the recent γ-ray flaring.”

4U 1755–33 M. S. E. Roberts and P. F. Michelson, Stanford University; L. R.

Cominsky, Sonoma State University; and F. E. Marshall, R. H. D. Corbet, and E. A. Smith, Goddard Space Flight Center, NASA, report: “On Jan. 30–31, the black-hole candidate 4U 1755–33 was observed by XTE to be in an off state for the first time. Preliminary analysis gives a conservative flux limit of < 1 mCrab, a factor of 100 below the source’s typical intensity. This offers an unprecedented opportunity to perform optical spectroscopic measurements of the companion star. The optical counterpart normally appears as a faint (V ∼ 18) blue object associated with reprocessed radiation from the accretion disk. Radial-velocity measurements over the 4.4-hr orbital period could lead to a determination of the mass function of the compact object. In addition, a detailed optical spectrum could determine the spectral type of the companion and resolve the issue of extreme underabundance of heavy metals in the accreted material.”

NOTICE TO SUBSCRIBERS Due to major recent configurations in the computer networking at the

Harvard-Smithsonian Center for Astrophysics, it is no longer possible to send these Circulars to e-mail subscribers via Decnet or Bitnet. Subscribers with such addresses should immediately supply us with a proper Internet address, to avoid delayed or lost Circulars.

1996 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green